OpenStudy (anonymous):

what is the decimal expansion of 4/9

OpenStudy (tkhunny):

1/9 = 0.11111111111111111111111111111... 2/9 = 0.22222222222222222222222222222... 3/9 = 0.33333333333333333333333333333... 4/9 = 0.444444444444444444444444444... I'm seeing a pattern.

OpenStudy (anonymous):

the answer choices r a. 0.49 _ b. 0.4 c. 0.4 d. 4.9

OpenStudy (anonymous):

B

OpenStudy (anonymous):

thx soooooooooooooo much